The below table presents the production of a company: L TP AP MP 0 0 -...

  1. The below table presents the production of a company:

L

TP

AP

MP

0

0

-

-

5

10

2

2

10

30

3

4

15

60

4

6

20

100

5

8

  1. Calculate the quantity that can be produced by 7 workers.
  2. How many workers needed to produce 48 units?
  3. Suppose that the company has 7 workers. If the company wants to increase the production by 30 units, how many more workers should hire?
Answer #1

a. Total product with 5 labor=10

MP between 5 th to 10th labor= 4

MPn=TPn-TPn-1

TP with 6 labor= 10+4=14

TP with 7 labor= 14+4= 18

Quantity produced by 7 workers= 18

b) 30 units are produced by 10 workers

MP between L= 30 & L= 60 is 6

So TP with 11 workers= 30+6=36

TP with 12 workers= 36+6=42

TP with 13 workers= 42+6=48

So 13 workers are required to profuce 48 units.

c) When 7 workers are there output is 18.

Increase the output by 30, 18+30= 48.

13 workers are required to produce output =48

13-7 =5

So 5 more workers are required.

If it helps kindly upvote